OpenStudy (anonymous):

PLEASE HELP!! Find the equation in standard form which passes through (–1, 4) and (0, 9)

OpenStudy (anonymous):

\[m=\frac{ y2-y1 }{ x2-x1 }\] \[y - y1 = m*(x - x1)\]

OpenStudy (anonymous):

The slope is 5 right?

OpenStudy (anonymous):

y-9 =5(x-0)

OpenStudy (anonymous):

yes, but still not done yet solve for y y = mx + b

OpenStudy (anonymous):

I don't really know how to use that equation/: i never learned it

OpenStudy (anonymous):

y-9 =5(x-0) distribute the 5 what do you get?

OpenStudy (anonymous):

y-9 = 5x - 0 ?

OpenStudy (anonymous):

yes now move the 9 to the other side what do you get?

OpenStudy (anonymous):

then subtract 5x from both sides?

OpenStudy (anonymous):

no you need to add 9 to both sides

OpenStudy (anonymous):

y-9 = 5x +9 +9 y = 5x + 9

OpenStudy (anonymous):

wait but I need to put it in standard form

OpenStudy (anonymous):

y = 5x + 9 ok subtract the 5x and you will be good

OpenStudy (anonymous):

Standard Form: Ax + By = C slope-intercept form: y = m x + b

OpenStudy (anonymous):

yeah i know but I think I got it now I figured it out thank you for helping me understand tho c: